OpenStudy (anonymous):

The sum of the ages of the brothers Joe and Paul is 49. Eleven years ago Joe was twice as old as Paul. How old is each now?

geerky42 (geerky42):

Let J be age of Joe. Let P be age of Paul. We are given that J+P=49 and J-11=2(P-11)

OpenStudy (anonymous):

Why would the 2 be with the p?

OpenStudy (anonymous):

wouldn't it be 2(j-11)=p-11 ?